LHC Startup

Runs

RUN Events Start Stop Beam Beam note BPTX BSC EG5 Trigger Note Note
122023 42763841 2009.11.21 11:55:56 2009.11.21 11:55:45 No            
122011 3290085 2009.11.21 11:19:39 2009.11.21 11:44:27            
122004 1622986 2009.11.21 10:54:21 2009.11.21 11:14:39            
121999 155 2009.11.21 08:38:01 2009.11.21 09:37:02            
121998 3 2009.11.21 07:50:20 2009.11.21 08:30:27            
121997 96 2009.11.21 06:58:02 2009.11.21 07:40:13            
121995 220 2009.11.21 05:26:01 2009.11.21 06:46:42            
121993 279 2009.11.21 03:52:51 2009.11.21 05:07:02 B1 Splashes miss   X    
121982 1 2009.11.21 01:50:16 2009.11.21 02:09:54 B2 Splashes X     1 splash All events to the ErrorStream
121981 4 2009.11.21 01:44:13 2009.11.21 01:47:52 X     4 splashes
121980 18 2009.11.21 01:20:33 2009.11.21 01:35:29 X     GCT Out, EG5 timed in at 51 for the SLBs
121976 27 2009.11.21 00:46:34 2009.11.21 01:11:37 Capture X    
121964 180679 2009.11.20 21:31:32 2009.11.20 23:35:17 B2 Splash+Circ.+Capt. X X   BSC added after splashes
and timed on the BPTX
~25000 turns at most
121962 89305 2009.11.20 20:45:02 2009.11.20 21:13:38 B1 Capture X X X  
Interrupted by trg when trying to apply the HTL prescale.
BGO Stop sent. DCCs Stopped.
121960 2132 2009.11.20 20:16:42 2009.11.20 20:40:50 Circulating X X X    
121949 1030 2009.11.20 19:15:13 2009.11.20 20:15:31 Fast track X X X Only 2 EG, Flooded by BSC  
121943 466 2009.11.20 17:21:42 2009.11.20 19:13:58 X X X  
121942 4488 2009.11.20 16:54:54 2009.11.20 17:14:54 X X X  

Friday

Saturday

Week 47

Friday

  • Super-stable until Monday
  • No collisions before end of the next week. 2 weeks estimated since circulating beam but lot of pressure to reach them earlier
  • Stability lock released on monday
  • It will be enforced again before the collisions

Report - Database Redundancy to be tested in the future - RCT Automatic Masking from FED vector - Daily planning - DCC fw rollback

Week 34

  • Magnet ramps down at 7.30 Tuesday morning, up at 11.00 Friday morning.

Notes

  • Wolfgang on holidays from Thursday on for 10 days.
  • Ken back on Monday
  • Sasha Singovski as replacement for 3 days
  • Mask CCU 32 in EE-5

Week 33

Monday

  • Call Pascal in the moning to figure out whether is realistic to test the EE+ trigger on Thurdays. If that is the case, check with Adi whether to take out of the global EE+ for 4/5 hours to hand it to Sasha for the HV scan.
  • SOAK light TOrbit Gap: starts at 3480, ends at 3564. The trigger gap ends at 3563 and lasts 35 bx backwards. Check where the SOAK light goes with the pfg, check what the EMTC people can do about the soak light triggers The trigger can extend the gap backwards for CRAFT if required. (expected between 3485-3495, two months ago )
  • DONE
  • ZS threshold in EE+ changed

Wednesday

  • HV scan, Wednesday starting between 8 am and 2pm.
    • Run stopped, EE+ out, laser out
    • EE(+/-) ramped down. HV settings changed, EE+/- ramped up.
    • EE- in again
    • EE+ to Sasha in the afternoon, out until next morning.
    • Sergey and Diogo alerted

  • Soak light
    • Trigger Masking
    • Software masking

Thursday playtime

Requests:
  • Pascal: EB trigger timing problem (1-2hrs), here are 2 board in EE- that are unplugged because the compact flash is corrupted (1hr), status of EE+ trigger (2hrs with GT-GCT-RCT). Total:~4-5hrs, out of which 2hr using RCT+GCT+GT.
  • ES: 1hr(LTC)+1hr(GCT)+1hr(firmware) = 3hrs
  • Jose': 2/3 hrs to debug the FOB problem

    • EE+ Trigger + SRP decisions [avra' bisogno di tutto EE+]
    • ES DCC firmware update tests
    • DCC weights with EB+/EB-
    • ES PVSS updates
    • ES DCC monitoring software updates

Friday

  • Friday Run Meeting
    • Data Quality Report, 20 minutes including the question time

Week 32

  • Led soak light (switched on by Sahsa w/o warning)at 1 kHz. It's not correctly timed in and it leaked into the physics events. -> RUN BAD
  • EB-5 , -6 noisy, the 2 EB- SMs affected by the TRK noise. -> Not a big deal. The DCC switches to the ZS and the ForceSuppression, FIFO Full flags are set. Nel DQM si vedono 5-100 eventi per RUN su due torri singoli. Su 40 Milioni
  • Saturday: Sistemati Ped Offsets. Pedestal Offset Scan.
  • Sasha HV scan: Ha bisogno di tutto EE+ [Togliere il laser dalla calib sequence: Cosmics-SR-No-Laser].
  • Readout: Tutte le DCC vanno bene in Global (non sempre il local). 2 casi di wrong event number, non si sa se sia un problema di dcc o myrinet. 3 DCC (EE+5,+6,+7) hanno mandato l'evento sbagliato, nello stesso crate. EE-1 venerdi' ha mandato l'evento sbagliato. (Sync Lost Draining DAQ status)
  • Cambiare la soglia del selective readout EE+ a 15 Lun Mattina. (soglia 1, la 2 non e' usata).

-- AlessandroThea - 18-Nov-2009

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2009-11-21 - AlessandroThea
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback